# A Review on Deep Learning Autoencoder in the Design of Next-Generation Communication Systems

### v1

@@ IX-C Nonlinear Activation Function
In summary, formula $N(2L-1)+L+F_{act}$ can be used to calculate the total number of FLOPs, where $F_{act}$ the FLOPs associated with the chosen activation function based on Table  V.
TABLE V: Number of FLOPs of Activation Functions
Activation Functions FLOPs
ReLU $N$
ELU $8N$
Leak ReLU $2N$
Tanh $13N$
SELU $9N$
Swish $8N$
Softmax $17N-$ 1
Sigmoid $7N$
X Literature Review of leveraging AE in Communication systems
Recent research on AE applications in communication systems focusses mainly on the following key research directions.
  • System-Level Design: focus on optimising communication systems holistically. Efforts in transceiver design aim to develop advanced architectures for efficient multiuser and varying channel scenarios. Complementing this is the push for end-to-end learning frameworks for complex scenarios, which unify traditionally separate processes such as channel coding, modulation, and equalisation. Research into high-performance code design explores capacity-approaching codes and innovative loss functions to improve power efficiency and communication reliability. Meanwhile, waveform optimisation leverages AEs to generate waveforms that meet constraints, ensuring high spectral efficiency and robust performance.
  • Channel and Signal Processing: Efforts address specific channel impairments and improve signal processing. Channel modelling mitigates issues such as fading and phase noise. DSP techniques improve equalisation and noise mitigation. Non-differentiable channels are tackled with GANs, reinforcement learning, and Kalman filters. Robust transmission ensures reliable performance in noisy and interference-prone environments.
  • Technology-Specific Applications: AEs are adapted for various communication technologies. MIMO systems use AEs for beamforming, CSI feedback, and interference management. OFDM systems address synchronisation, PAPR, and waveform optimisation. CRNs benefit from improved spectrum sensing and interference mitigation. Optical systems are used to deal with turbulence and absorption, while quantum communication employs hybrid AEs to handle noise and improve performance.
  • Functional Improvements: It focusses on redefining core communication tasks. Channel coding enhances reliability in complex scenarios. Modulation and demodulation adapt jointly to impairments. Semantic communication prioritises meaning over raw data, improving efficiency and handling semantic noise.
  • Computational and Practical Constraints: It focusses on resource efficiency and real-world applications. Resource-constrained systems optimise AE performance for hardware such as FPGAs. Hybrid learning strategies combine traditional methods with deep learning. AE interpretability improves transparency and scalability under varying conditions.
  • Emerging Applications and Novel Methods: AEs are applied to novel domains. Specialized applications include radar systems, underwater communication, and RIS networks, focussing on beamforming and signal reconstruction. New methods improve BER, spectral efficiency, and resilience in challenging environments.
AE-based Communication SystemsResearch issueApplication areasChannel modelAE approachWireless CommunicationSemantic CommunicationQuantum CommunicationOptical CommunicationAEArchitectureOptimisationApproachLayer type, ActivationLoss functionSystemDesignChannel/SignalProcessingTechnology-Specific App.Functional ImprovementsComputational &Practical ConstraintsMIMOAWGNRayleighOpticalVLCOWCUOCMSECEBCECustomfunctionDense, ResNetReLU, ELUBit-wiseSymbol-wiseAEVAE
Figure 8: Thematic Taxonomy of AE-based Communication Systems
Fig. 8 shows a thematic taxonomy for research related to AE in communication systems. It presents a hierarchical structure, starting from categorising various research issues, channel type modelled, AE architecture used in the current literature, including type of layers, loss functions, and application areas where the model has been applied, like Wireless Communication (WC), Optical Communication (OC) including Visible Light Communication (VLC), Optical Wireless Communication (OWC), and Underwater Optical Communication (UOC).
X-A Wireless Communication
X-A1 Transceiver Design
This section explores various aspects of AE technologies-enhanced transceiver design systems. Notable studies focus on handling channel variations, optimising constellations under specific channel models, and improving performance in complex environments such as Rician flat-fading channels, nonlinearities, multipath fading, and MIMO broadcast interference channels. Key studies highlight approaches for waveform optimisation under constraints and the use of advanced neural network designs, such as ResNet blocks, to enhance system performance under varied channel conditions.
MIMO Systems
Numerous scholarly articles have examined the application of an AE-based MIMO communication system for transceiver design. A system designed to handle random variations in the channel, particularly focussing on MIMO systems, multi-dimensional constellations (MDC), and Rician flat fading channels, is proposed in [68]. Dorner et al. [69] introduced a bit-wise AE that can achieve shaping gains in MIMO systems by considering the Gaussian channel model with ergodic Rayleigh channel matrices. It deals with the curse of dimensionality when scaling the architecture for larger bit configurations (beyond 8 bits) and the need for constellation optimisation tailored to specific receiver algorithms to enhance performance. Optimisation of communication reliability in reconfigurable intelligent surface (RIS)-assisted MIMO systems is explored in [70]. It replaced the transceiver and the RIS model with three fully-connected neural network (FCNN) models and jointly trained the proposed system to minimise BER. Furthermore, Song et al. [71] analysed how AE for interference channels learns to avoid interference in a rotated reference frame, with AWGN, open-loop MIMO, closed-loop MIMO, and MIMO broadcast interference channels.
OFDM Systems
An AE-based approach to optimise OFDM systems for transceiver design. Felix et al. [72] examined the ability of the AE-based OFDM system to handle nonlinearities and synchronisation issues without explicit compensation algorithms. It used a WiFi-inspired stochastic channel model derived from a Proakis type C tap-delay-line model, with AWGN and multipath fading. Aoudia et al. [73] advanced the joint design of transmit and receive, with constellation geometry and bit labelling to eliminate orthogonal pilots in scenarios with frequency and time-selective fading channels. The proposed AE used ResNet blocks with separable convolutional layers and the ReLU activation function, along with a separable two-dimensional (2D) convolution output layer. The design of OFDM waveforms that meet specific constraints on PAPR and ACLR has been addressed in [24, 41, 25]. The proposed AE model used ResNet blocks, which in turn used separable convolutional layers and a convolutional output layer. The BCE loss function designed to maximise achievable information rates while satisfying PAPR and ACLR constraints is suggested for designing OFDM waveforms. The first two papers [24, 41] employed ResNet blocks with ReLU activation, while the third paper [25] substituted ReLU with ELU to address the problem of vanishing gradients. In the same direction, Marasinghe et al. [74] addressed the key issue of Constellation Shaping under phase noise (PN) impairment for Sub-THz Communications. The study used the AWGN channel model, which simplifies the analysis by assuming perfect channel equalisation while focussing on PN impairment. The loss function is formulated in an augmented Lagrangian form, which is used to optimise the constellation design robust to PN while maintaining low PAPR.
@@ X-A1 Transceiver Design
TABLE VI: Comparison of AE-based models for Transceiver Design
Ref. Main Contribution Model(s) Used Dataset/Data type/Channel model Best performance value Limitation
[25] PAPR, ACLR reduction ResNet blocks (sep. conv, ELU) Synthetic, AWGN 8.5 dB PAPR reduction Constraint evaluation complexity
[27] Non-differentiable channel Tx: Dense (ELU), Rx: Dense (ReLU), RBF Synthetic, AWGN, RBF Comparable to supervised Slower convergence on AWGN
[41] Joint Tx/Rx, constellation, labelling ResNet blocks (sep. conv, ReLU) Simulated, AWGN, multipath ACLR --29.95 dB, PAPR 5.60 dB NN architecture complexity
[68] Adapt to random fading Symbol differentiation, dense Rician fading, AWGN Superior SER in high SNR Optimised for specific channel
@@
[69] Curse of dimensionality Bit-wise AE, dense (ReLU) Complex Gaussian APP receiver for k=8 Scalability limitations for k¿8

### v2

@@ V-G3 Nonlinear Activation Function
In summary, formula $N(2L-1)+L+F_{act}$ can be used to calculate the total number of FLOPs, where $F_{act}$ the FLOPs associated with the chosen activation function based on Table  IV.
TABLE IV: Number of FLOPs of Activation Functions
Activation Functions FLOPs Activation Functions FLOPs
ReLU NN ELU $8N$
Leak ReLU $2N$ Tanh $13N$
SELU $9N$ Swish $8N$
Softmax 7N17N-1 Sigmoid $7N$
VI AE-Based Solutions to Domain-Specific Challenges in Communication Systems and their Limitations
This section provides a detailed analysis of communication challenges and AE roles in various domains, including wireless, optical fiber, free-space optical, visible light, semantic, quantum and underwater communications. First, it highlights the general limitations and challenges in using AE in communication systems, Then for each area, it outlines key difficulties within each area, such as complexity, channel variability, interference, data and training issues for machine learning, system and hardware limitations, security, performance, and interpretability, while also discussing potential solutions of using deep learning AE architectures as promising approaches. Moreover, it also explores various limitations encountered when implementing deep learning AEs in these specific areas of communication systems.
VI-A Wireless Communication
Wireless communication systems face a wide array of challenges, spanning from fundamental issues like channel impairments and interference to specific challenges in wireless sub-field systems such as multiple-input multiple-output (MIMO), OFDM, integrated sensing and communication (ISAC), and CR technology. Figure 6 shows an overview of typical challenges of wireless communication systems, where the problem is addressed by AE-based architecture.
Wireless CommunicationsSolutionsChallenges Channel Impairments and Variability: lack of exact channel models, mismatch between model and real conditions, noise/quantization robustness, channel coding/equalization Interference Management: dynamic and unknown interference, co-channel effects, trade-off between suppression and performance MIMO Systems: curse of dimensionality, decentralized processing schemes, signal detection, nonlinear/interference distortion, CSI-efficient feedback, ZIC OFDM Systems: inherent waveform characteristics, PAPR, ACLR, synchronization, phase noise ISAC Systems: JCAS, difficulties in detecting multiple radar targets, balance sensing and communication Model-free AEs, stochastic approximation algorithms, GAN-based channel modeling, Turbo AE, VAE equalizer, denoising AE Interference-aware training, joint transmitter–receiver AE design, adaptive deep learning Bit-wise AE, BWAE-MIMO, SWAE-MIMO, VAE-based, AE-based architecture, PRVNet AE-driven waveform design methodologies, AE-based GCS, MC-AE AE-based Joint constellation shaping, AE-based for permutation-invariant encoding
Figure 6: Wireless communication challenges and solutions
VI-A1 Communication System Design
Table V lists the recent literature that focuses on the design of AE-based communication systems. Real-world conditions introduce significant channel impairments that degrade signal quality in wireless communication systems. These include extreme variability such as noise and distortion, which complicate the development of comprehensive and accurate channel models [61]. A central challenge in this context is the unavailability of an exact channel model, which hinders the optimization of the AE parameters [38, 61]. Moreover, traditional modeling approaches often fail to capture complex relationships in channel behavior and are inadequate to accurately represent real world conditions [32]. Training systems with mismatched channel models can also lead to significant performance degradation [62].
TABLE V: State-of-the-art comparison of AE-based models in system design
[38] Proposes stochastic approximation for AE training without exact channel model Stochastic Approximation AE OFDM, Fading Channels Maintains robustness under channel uncertainty Complexity in implementation
[32] Uses GANs for real-time channel modeling with AEs GAN-based AE Real-world channels Effective channel representation and interference handling Needs high-quality training data
[47] End-to-end AE-based full communication system End-to-End AE Single-antenna, impaired channels Adaptable, all-in-one AE framework No direct comparison to baselines
[61] Compares model-assumed vs model-free channel AEs GAN-CAE, RL-CAE AWGN, Rayleigh Channels Comparable BLER, robust under complex channels Model-free requires more training
[62] Highlights AE performance loss from channel-model mismatch Trainable Channel AE Synthetic channels Insight into model mismatch sensitivity Lack of practical mitigation strategy
[63] Evaluates model-free AEs under various channel models Dense AE AWGN, Rayleigh Performance stable under different impairments Training time increases with channel complexity
[64] Introduces end-to-end AE without differentiable channel model Dense AE AWGN, RBF, fiber-optic Comparable to model-based AE, hardware-implementable Assumes accurate feedback for training
[65] Develops interference-aware AE with randomized smoothing Interference-aware AE Dynamic interference, jamming Robust to unknown interference, improved fidelity Complexity in balancing suppression and accuracy
[66] Highlights AE difficulty in adapting to jamming and dynamic interference End-to-End AE Adversarial channel conditions Realistic interference mitigation examples No general framework proposed
[30] Adaptive AE for dynamic interference management in beyond-5G systems ADL AE Multi-user Gaussian interference channels 2 dB improvement in BER over conventional schemes Performance highly dependent on interference estimation and higher complexity
[67] Proposes TurboAE with near-optimal AWGN performance TurboAE AWGN, Non-AWGN Exceeds LDPC/Turbo in mid-range block length High SNR training requires large memory
[68] Integrates modulation with TurboAE (TurboAE-MOD) TurboAE-MOD Non-Gaussian channels More robust than traditional modulation Training requires differentiable channels
[69] AE-based Turbo coding for 1-bit quantized systems ConvAE + Turbo Code AWGN, Quantized Inputs Enables 16-QAM, lowest BER for 1-bit RX Suboptimal AE training leads to local minima
[…]
[71] VAE equalizer for blind decoding w/o pilots VAE Equalizer Linear/Nonlinear ISI BER $10^{-3}$ at 6 dB, pilot-free High sensitivity to hyperparameters
[72] Bi-GRU AE optimized for ISI channels Bi-GRU AE ISI, AWGN Better than LDPC at low SNR Needs performance boost at high SNR
[73] VAE-based JSCC for image and Gaussian sources VAE with MoE Gaussian, Laplace, Image Data SDR near Shannon limit, robust to mismatch JSCC lacks modularity and standards
[74] Mutual info-based AE loss for capacity-approaching codes MI-Regularized AE Unknown Channels Better mutual info and BLER vs cross-entropy Not scalable to high dimensions
[75] AE + DAE for robust latent image transmission Stacked AE/DAE Latent Vectors (Image) Improved CA, MSE resilience to quantization Assumes known latent vector, ideal conditions
[76] LSTM DAE for modulation classification on edge devices LSTM Denoising AE RadioML2018.01A, PSD Data 96.56% accuracy, real-time on Raspberry Pi 4 Limited advanced modulation coverage
[77] AE with non-negativity for improved mod classification AE + Constraint 4th-order cumulants 100% accuracy at 10-15 dB SNR Limited to low-dim features
[78] Stacked DAE for modulation classification under noise Stacked Denoising AE Synthetic modulated signals Higher accuracy than conventional classifiers Lacks real-world validation
[79] Low-complexity AE for coded comms at low SNR Low-Complexity AE AWGN Better BER than 16-QAM in low SNR Needs validation vs modern codes
[80] Early low-complexity AE showing BER gains Dense (ReLU) Binary, AWGN Beats 16-QAM in ideal cases Idealized assumptions only
Therefore, AEs have been explored for designing entire communication systems. Morocho et al. [47] showcased the versatility of AEs by developing an end-to-end communication system that replaced traditional transmitter and receiver tasks. Their approach optimized system operations and was effectively adapted to channel impairments. However, the core challenge being addressed is the sub-optimality of conventional block-based designs. To address the challenge of unavailability of an exact channel model, AEs can be designed to dynamically adjust to measured channel characteristics, both model-assumed and model-free channel AEs have been proposed as viable methods for mitigating channel impairments and improving performance [63, 61, 64]. A recent innovation involves an end-to-end learning algorithm based on AE that eliminates the need for a differentiable channel model, yet achieves performance comparable to traditional model-based training across various channel conditions [64]. Furthermore, stochastic approximation algorithms offer a solution to the absence of exact channel models by enabling robust learning under uncertainty [38]. Additionally, deep learning AE architectures that use generative adversarial networks (GANs), facilitate real-time channel modeling and adaptation [32]. Both model-assumed and model-free channel AEs are primarily presented as solutions designed to address channel impairments in communication systems. However, when comparing the two approaches, some characteristics could be seen as relative limitations. First, model-assumed AEs are noted to outperform model-free AEs in simpler environments [61, 64]. This implies that model-free AEs might have limitations in achieving peak performance when the channel conditions are well-behaved and can be accurately modeled. Second, model-free AEs require more training time compared to model-assumed AEs, although they can achieve comparable performance. This longer training time is a limitation in terms of efficiency and practical deployment speed.
Interference presents a major challenge in wireless communication systems, particularly in densely populated spectral environments, where isolating the desired signal becomes increasingly difficult. In multi-user multiple access channels (MAC), inter-user interference further complicates reliable communication, which requires robust interference management strategies. End-to-end communication frameworks that use AEs inherently capture a wide range of impairments, including interference, through joint transmitter–receiver optimization [32]. Moreover, AE-based systems that incorporate interference-aware training techniques, such as randomized smoothing demonstrate a strong ability to suppress unknown and dynamically varying interference sources, thereby improving resilience and communication fidelity [65]. Adapting AE-based communication systems to dynamic and diverse interference scenarios is non-trivial, as achieving effective suppression without compromising system performance remains a significant hurdle. Several open challenges and limitations are identified when using an AE-based system to address interference impairments. First, adapting AE communications to various interference scenarios, including co-channel interference, is considered complex [65, 66]. This complexity extends to handling unknown and dynamic interference conditions, such as jamming effects. Second, achieving effective interference suppression while maintaining communication performance is highlighted as a significant challenge when using AEs in the presence of interference [65]. There is also the challenge of balancing accuracy and robustness in these conditions. Lastly, there is a need for modeling interference in addition to channel effects. This implies that simply capturing the channel may not inherently handle complex interference scenarios without specific design considerations.
Wu et al. [30] proposed a novel adaptive deep learning (ADL) based AE to address the challenge of dynamic interference in multi-user Gaussian interference channels, which is particularly challenging under high interference conditions where standard constellations are not optimal. It demonstrated significant performance improvements over conventional n-psk and n-QAM modulation schemes under high interference conditions. However, it considered adapting modulation and decoding strategies to specific channel conditions and interference levels.
The importance of AEs for channel coding and decoding comes from their potential to redefine communication system design through end-to-end optimization. Therefore, Jiang et al. [67] proposed Turbo AE (TurboAE), which achieved state-of-the-art performance in conventional AWGN channels for moderate block lengths and exceeded traditional codes such as the LDPC and Turbo codes. The design of TurboAE automates channel coding through deep learning using novel training algorithms and structures for the encoder and decoder. However, a limitation noted is that performance at larger block lengths requires extensive training memory, which remains challenging due to the rarity of error events. Jiang et al. [68] also introduced TurboAE-MOD, which combines channel coding and modulation within an end-to-end AE framework. This approach exhibited enhanced reliability and robustness, especially under non-Gaussian channel conditions, compared to conventional modulation techniques. The challenges in this approach include the requirement for differentiable channel models for backpropagation during training, although model-free learning algorithms that do not require differentiable channels have also been proposed and implemented on hardware [64, 61]. Furthermore, Balevi et al. [69] proposed a system combining turbo codes with AE for one-bit quantized receivers. This deep learning-based coding scheme was designed for AWGN channels under one-bit quantization constraints with the aim of enhancing error correction performance. The proposed method outperformed conventional turbo codes and enabled operational one-bit quantization for higher modulation orders like 16-QAM. However, the research also acknowledges challenges such as the unattainability of perfect AE training, which leads to suboptimal training, and the inherent difficulty in training deep networks, often resulting in local minima rather than global optima.
Van et al. [70] introduced a single user non-coherent energy AE (NC-EA) system, which uses DNN for encoding and decoding. This system operates by allowing the decoder to use energy combined from all receive antennas, efficiently recovering transmitted data without requiring CSI estimation. Novel DNN structures were developed for both uplink and downlink NC-EA multiple access schemes. The simulation results demonstrated the superiority of the proposed NC-EA schemes over baseline methods in several key performance indicators, including reliability, flexibility, and complexity. The NC-EA system’s ability to avoid CSI estimation simplifies the receiver design, which can be particularly attractive for applications requiring reliable, low latency, and low complexity connectivity, such as machine-type communications (MTCs). While It offers significant advantages, the need for effective strategies to handle the mismatch between training models and real-world channel conditions is imperative.
Another application of AEs is channel equalization. Caciularu et al. [71] introduced VAE equalizers specifically for the purpose of blind equalization. It aims to reconstruct transmitted symbols with unknown channel parameters, including impulse response and noise variance, without relying on pilot symbols. VAE equalizers have shown significant improvements in error rates for reconstructed symbols compared to existing blind equalization methods and enable faster channel acquisition. Critical analysis notes that VAE equalizer performance can be significantly influenced by hyperparameters, and while effective for linear impairments, extending them to handle nonlinearities remains an area for future research. Zhang et al. [72] introduced a Bi-GRU-based AE tailored for ISI channels that offers flexible coding rates and outperforms convolutional and LDPC codes in low SNR regimes. Nevertheless it requires further enhancements at high SNR.
AE have also been applied to joint source-channel coding (JSCC), where the goal is to encode the source data directly into symbols for transmission over a noisy channel to optimize for end-to-end distortion. Saidutta et al. [73] developed a VAE-based JSCC scheme, which achieved near-optimal SDR for Gaussian sources while maintaining robustness across different channel conditions. The approach also demonstrated superior performance on image datasets compared to traditional methods. Saidutta et al. [73] developed a VAE-based JSCC scheme, which demonstrated effective performance over analog noise channels and achieved performance comparable to or better than existing methods for Gaussian sources over AWGN channels. A key finding is that this methodology generalizes to other source distributions, including Laplace and image datasets, which shows consistent improvements and demonstrating robustness to changes in channel conditions. Despite these successes, JSCC faces challenges such as residual noise accumulation over multiple hops, a lack of coding standards for learned methods, loss of modularity that requires separate designs for different source types, and issues with non-differentiable components in standard codecs that hinder training.
Another significant area is the design of codes that approach theoretical channel capacity using novel loss functions. Letizia et al. [74] proposed an AE loss function that incorporates mutual information to design capacity-approaching codes. This loss function is inspired by the information bottleneck method and channel capacity. The aim was to design capacity-approaching codes under power constraints without prior channel knowledge by jointly maximizing mutual information and minimizing cross-entropy. It showed improved decoding schemes and rates and that traditional cross-entropy loss functions can lead to overfitting and suboptimal signal representation. The MI regularization term can help to control the information stored in the latent representation. However, this research faced critical challenges that the AE struggled to scale with high code dimensions, which risk local minima or failure to converge to zero-error codes. Furthermore, the mutual information estimation methods used, such as the MINE block, can be unreliable with a large number of samples. The experiments carried out were limited to a 3-bit system. This suggested a need for further experimentation and hyperparameter tuning for in-depth analysis and generalization to higher bit scenarios.
Robustness against noise and quantization is crucial for practical systems. Khan et al. [75] combined basic AE and denoising AE to improve robustness in transmitting latent image representations under noise and quantization, which demonstrated significant improvement in image reconstruction quality and classification accuracy (CA). Nonuniform quantization of the parameters of the AE was found to significantly reduce MSE and improve CA, which closely matches the performance without quantization. Conversely, fewer quantization bits led to increased MSE and decreased CA. Critical analysis notes the vulnerability of digital transmission using conventional AEs to quantization and channel noise. Limitations included scenarios in which the original transmitted latent vectors were assumed available, neglecting real impairments, and disregarding the effects of digitization and quantization of latent variables and parameters in practical systems.
In the realm of modulation classification, Ke et al. [76] designed a LSTM-based denoising AE for radio technology and modulation classification. It achieved high classification accuracy, reaching 96.56% at high SNR on the RadioML2018.01A dataset. A key contribution was the development of a compact neural network architecture, specifically designed for real-time classification on low-cost computational platforms like the Raspberry Pi 4. The proposed LSTM Denoising AE aimed to enhance classification accuracy while simultaneously reducing computational complexity. It was noted that existing spectrum monitoring techniques often struggle with achieving the required accuracy and computational efficiency on low-cost platforms. Similarly, Ali et al. [77] introduced an AE with non-negative constraints, which achieved 100% classification accuracy at high SNR (10-15 dB) and demonstrated robustness under fading conditions. This approach outperformed traditional sparse AEs in both reconstruction error and classification accuracy. Furthermore, Zhu et al. [78] presented a stacked denoising AE for rapid and accurate modulation classification. The proposed approach outperforms traditional methods, especially in noisy environments. For the classification task, despite achieving high accuracy and efficiency on low-cost platforms. The difficulty in distinguishing certain modulation schemes with similar spectral properties persists.
Researchers have focused on developing low-complexity AE architectures for communication systems. Rajapaksha et al. [80, 79] developed a low-complexity AE that aimed to reduce model dimensionality and training complexity. It outperformed convolutional coding, particularly within the range of 0-5 dB $E_{b}/N_{0}$ in the low to moderate SNR ranges and also achieved better BER than the 16-QAM half rate with hard decision decoding. However, this performance analysis was based on an ideal communication system assuming perfect synchronization, and further research is needed to evaluate its performance in non-ideal conditions and against more advanced channel codes like LDPC and polar codes.
VI-A2 MIMO Systems
@@
Table VI lists the state-of-the-art comparison of AE-based models for MIMO systems.
TABLE VI: State-of-the-art comparison of AE-based models for MIMO systems
@@
[81] Curse of dimensionality Bit-wise AE, dense (ReLU) Complex Gaussian APP receiver for k=8 Scalability limitations for $k>8$
@@
[89] Efficient CSI feedback with robustness to noise PRVNet (VAE-based) Indoor and outdoor FDD-MIMO scenarios NMSE: 27.7-27.7 dB (indoor), 13.9-13.9 dB (outdoor) at CR=1/4 Limited to feedback channels with specific compression ratios
[90] Joint quantization and compression for CSI feedback AE with entropy bottleneck layer Indoor (5.3 GHz) and outdoor (300 MHz) CSI feedback NMSE: 24-24 dB (indoor), 13-13 dB (outdoor) at BPP=3.0 Requires specific tuning for optimal entropy bottleneck performance
The integration of deep learning AE architectures into MIMO systems introduces significant challenges, primarily due to the curse of dimensionality. As the number of antennas and the bit representation in vector constellations increases, AE architectures experience exponential growth in complexity, which leads to computational and training difficulties. To address this issue, Dörner et al. [81] introduced a bit-wise AE designed to achieve shaping gains in MIMO systems introduced a bit-wise AE. However, the system is limited to constellations with up to 8 bits due to training difficulties for larger configurations. Complexity increases exponentially with setups exceeding 8 bits, which led to decreased mutual information performance. This curse of dimensionality presents a significant barrier for AE architecture, especially with binary detection and decoding.
Expanding on AE-based approaches for signal detection, Bui et al. [82] introduced two innovative systems, BWAE-MIMO and SWAE-MIMO, designed to leverage AEs for improved signal detection performance in MIMO systems. Notably, the BWAE-MIMO system, equipped with a radio transformation network (RTN) detector, demonstrated BER performance comparable to the maximum likelihood detection (MLD) approach when evaluated in 2x2 MIMO configurations. However, several critical considerations temper the practical application of AE-based MIMO systems. Scaling AE architectures to accommodate larger numbers of bits per vector symbol presents significant challenges due to the curse of dimensionality and associated training difficulties. Furthermore, implementing AE systems in real-world scenarios is complicated by the requirement for differentiable channel models (or effective workarounds) and the persistent challenge of ensuring robustness and adaptability across varying channel conditions not perfectly matched during training.
In a similar vein, Zhao et al. [83] developed a VAE-based signal detection architecture, specifically adapted to MIMO-OFDM-IM systems. It achieved BER performance that was not only close to existing detection algorithms but also featured reduced runtime, thus underscored the efficiency of deep learning-driven solutions. Nevertheless, this approach also faces challenges. While VAEs have shown promise in areas such as blind channel equalization and handling noisy latent codes, their performance can degrade in high-resolution or high-dimensional scenarios. Achieving optimal results often requires meticulous hyperparameter tuning, and ensuring robustness against complex noise scenarios remains an area requiring further study.
In the domain of millimetre-wave (mmWave) communications, Tao et al. [84] introduced an innovative AE-based hybrid beamforming method (AE-HB). By effectively mapping the traditionally non-convex hybrid beamforming optimization problem onto an AE neural network, their approach transformed it into a more tractable neural network training process. This method demonstrated a significant 3 dB improvement in BER compared to conventional linear beamforming systems. However, AE-based transceivers trained on specific channel models can experience performance degradation when there are discrepancies between the training model and the actual transmission link. Furthermore, end-to-end AE systems often require differentiated channel models for backpropagation, which can hinder development unless workarounds are used.
On the other hand, Cherif et al. [85] developed an AE end-to-end learning approach aimed at mitigating nonlinear distortions, such as those introduced by high-power amplifiers in massive MIMO systems. The proposed approach achieved near-linear SER performance at high SNRs with 16-QAM modulation, which significantly outperformed traditional distortion mitigation techniques. However, achieving optimal performance often requires a careful selection of network parameters and hyperparameters, which can be a complex process. Learned constellations optimized for specific receiver algorithms may not achieve optimal performance with different receiver types. In addition, generalizability across different types and sources of nonlinearity needs thorough investigation. Massive MIMO systems present another layer of complexity [93], particularly with decentralized processing schemes. These systems are constrained by interconnect bandwidth, and current compression techniques often fail to reduce data transfer rates to acceptable levels. As a result, overall efficiency and scalability are significantly impaired. This motivates the need for innovative solutions to mitigate data throughput demands. Sahay et al. [86] investigated the vulnerability of DNNs used for power allocation in massive MIMO networks to adversarial attacks. They developed a DAE specifically for mitigating adversarial attacks in massive MIMO power allocation networks. It effectively filters adversarial perturbations before inputting data into the deep neural networks for power allocation. The research demonstrated the DAE’s effectiveness without requiring retraining of the AE deep learning models and showed that it outperforms existing benchmarks for regression-based adversarial attack mitigation. Crucially, the DAE retains accurate performance in the absence of attacks and operates with low computational overhead, while addressing the high computational overhead identified as a limitation in traditional optimization-based methods.
Optimization of communication reliability in reconfigurable intelligent surface (RIS)-assisted MIMO systems is explored in [87]. It replaced the transceiver and the RIS model with three fully-connected neural network (FCNN) models and jointly trained the proposed system to minimize BER. The findings indicated significant gains in BER performance compared to benchmarks and demonstrated robustness even when perfect channel information is unavailable. Critically, it is clear that the need for a framework robust to imperfect channel information and the increased complexity with efficient signal detection/processing . Furthermore, Song et al. [17] analyzed how AE for interference channels learns to avoid interference in a rotated reference frame, with AWGN, open-loop MIMO, closed-loop MIMO, and MIMO broadcast interference channels. The study demonstrated that end-to-end AE learning can outperform traditional human-engineered transceivers in memoryless MIMO and MU systems without requiring prior knowledge of communication-theoretic principles. However, the study identifies potential pitfalls in interpreting learned communication schemes, particularly regarding interference management and the AE learning of interference avoidance in a rotated reference frame. Additionally, the AEs considered have a fixed data rate, which limits their adaptability for rate adaptation, and the neural network parameters used may not be fully optimal.
Jang et al. [88], Hussien et al. [89], and Ravula et al. [90] each addressed the challenge of CSI-efficient feedback in MIMO systems through innovative deep learning approaches. The first proposed a deep AE-based method that significantly reduced normalized MSE and demonstrated robust performance under feedback errors and delays, which ensured reliable CSI reconstruction. The latter introduced PRVNet, a VAE-based model optimized for high-compression scenarios. PRVNet achieved exceptional NMSE values of 27.7-27.7 dB in indoor environments and 13.9-13.9 dB in outdoor settings in a compression ratio of 1/4, which outperformed many existing methods. Similarly, the latter proposed an entropy bottleneck layer within an AE-based feedback framework, which enabled joint optimization of bit rate and distortion. This approach achieved NMSE values of 24-24 dB for indoor channels and 13-13 dB for outdoor channels at specific bit-per-pixel (BPP) values, which demonstrated superior reconstruction quality over traditional methods. Collectively, these studies highlighted the potential of AE and VAE-based techniques to improve CSI feedback efficiency and accuracy in diverse wireless environments. However, assuming that the CSI information is available at the transmitter may not always be the case. Furthermore, more studies are required on the impact of AE-based compression on CSI accuracy.
Furthermore, Zhang et al. [91] proposed a deep AE-based architecture for Z interference channels. They proposed system introduced interference-aware constellations that adapt their shape based on interference intensity with the explicit goal of minimizing the BER. They addressed the challenge of the two-user Z-interference channel (ZIC) with finite-alphabet inputs, and the conventional finite-alphabet approaches in this context are their reliance on predefined constellations that do not adapt to interference. The structure was also extended to work effectively with imperfect CSI, directly addressing the challenges posed by estimation and quantization errors. A notable feature of their design was the inclusion of an I/Q power allocation layer, which enables the generation of nonuniform constellation shapes, thereby enhancing robustness against interference. The simulation results indicated performance improvements of up to two orders of magnitude compared to standard methods. In another innovative effort, Shin et al. [92] introduced the AE framework based on amplitude phase quantization with neural networks (APQNN) for relay systems, which achieved significant BER improvements. Their approach demonstrated performance close to $10^{-3}$ BER at high SNR levels, which emphasized its effectiveness in optimizing the relay system.
TABLE VII: State-of-the-art comparison of AE-based models for OFDM systems
[53] optimize discrete distributions Tx: Dense (ReLU, linear), Rx: Dense (ReLU) AWGN, Rayleigh fading Close to capacity on AWGN Finding optimal marginal distribution is challenging
[49] Joint Tx/Rx, constellation, labelling ResNet blocks (sep. conv, ReLU) Simulated, AWGN, multipath ACLR -29.95 dB, PAPR 5.60 dB NN architecture complexity